OpenStudy (anonymous):

Four friends went scuba diving. According to the following statements, who spotted sea life at the fastest rate? Lance saw 2 reef sharks in 6 minutes. Jasmine saw 3 sea cucumbers in 8 minutes. Robert saw 4 jellyfish in 15 minutes. Shilpa saw 3 pencil urchins in 10 minutes.

OpenStudy (anonymous):

All you need to do is divide each one to see how fast they all spotted sea life in a minute. 2/6 = 1/3mins 3/8 = 1/2.7 mins 4/15 = 1/3.75 mins 3/10 = 1/3.3 mins Answer is Jasmine.